For the a Poisson random variable x generating function G(t)= exp^lambda(t-1) show that the expected value E(x) and variance (x) are given by lambda and lambda respectively.


1
Expert's answer
2022-01-14T09:44:59-0500

"G(t)=e^{\lambda(e^t-1)}"

"E(x)=G'(0)"

"G'(t)=\lambda e^te^{\lambda(e^t-1)}=\lambda e^{\lambda t(e^t-1)}"

"G'(0)==\lambda e^{\lambda 0(e^t-1)}=\lambda"


"var(x)=E(x^2)-[E(x)]^2"

"E(X^2)=G''(0)"

"G''(t)=\lambda e^{\lambda e^t+t-\lambda}(\lambda e^t+1)"

"G''(0)=\lambda e^{\lambda e^0+0-\lambda}(\lambda e^0+1)=\lambda(\lambda+1)=\lambda^2+\lambda"

From the first part, "E(x)=\lambda" . Thus;

"var(x)=E(x^2)-[E(x)]^2=\lambda^2+\lambda-\lambda^2=\lambda"